OpenStudy (anonymous):

Will give medal! Find the slope of the line through the pair of points. A(2, –3), P(2, 9)

OpenStudy (anonymous):

You have to use the slope intercept formula y2-y1 / x2-x1 = slope

OpenStudy (anonymous):

9 + 3 / 2 -2 = 11/0 You are only moving 11 spaces, there is no definite slope. Does that make sense?

OpenStudy (anonymous):

It does thank you very much!